the force between two charge 2uC and 4uC is 24N when they are separated by a certain distance in free space

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
Force = K * q1 *q2 /r²
 
24 = 9 * 10^9 * 4 *2 * 10^-12 /r²
 
r² = 3 * 10^-3
 
r² = 30* 10^-4
 
r = 5.45 *10^-2 metres
 
r = 5.45 cm